OpenStudy (anonymous):

Find the equation of the line that passes through (4,3) and is parallel to the line 3x-y=7 plz help

OpenStudy (anonymous):

so first you want to figure out the slope. if two lines are parallel, it means they have the same slope. can you tell me what the slope is of 3x-y=7?

OpenStudy (anonymous):

im looking at the textbook and it doest show me how to find the slope

OpenStudy (anonymous):

once you have the slope, you want to plug the two points (4,3) into the point-slope form: Y-Y1 = m (x - x1) where m is the slope and x1 and y1 represent the given point

OpenStudy (anonymous):

do you know how to put 3x-y=7 into slope intercept form, where y=mx+b where m = slope, and b = the y intercept?

OpenStudy (anonymous):

no i dont can u plz show me?

OpenStudy (anonymous):

of course ^.^ so basically you want to make 3x-y=7 look like y=mx+b

OpenStudy (anonymous):

so what do you think would be the first step, to try to get that y alone?

OpenStudy (anonymous):

ok so im guessing b is the y-intercept right?

OpenStudy (anonymous):

yup, but you want to get that y alone to figure out the slope

OpenStudy (anonymous):

i have no clue im srry

OpenStudy (anonymous):

no worries so you have this equation 3x-y=7, and you want to get the y alone - basically move everything from one side of the equation, to the other, until y is by itself.. first thing you want to do is move that 3x over, and you can do that by subtracting it from both sides if you do that, you're going to cancel it out on the left side, and "add it in" on the right side.. does that make sense? so you're going to have -y=-3x+7

OpenStudy (anonymous):

now that looks a lot more like y=mx+b, right? but we don't want the y to be negative.. what do you think we should do to make it positive?

OpenStudy (anonymous):

ok u have to take out the 3x?

OpenStudy (anonymous):

yeah, that's the first step, and then you want to make the y positive

OpenStudy (anonymous):

make the y positive by i dont know how

OpenStudy (anonymous):

you can do that by multiplying the y by -1, and whatever you do to one side of the equation, you have to do to the other.. so this is what it looks like: (-1)y=(-3x+7)(-1) or (-1)(y)=(-1)(-3x)+(-1)(7)) does that make sense?

OpenStudy (anonymous):

can you tell what the final equation is?

OpenStudy (anonymous):

oops (-1)(-y) ^.^

OpenStudy (anonymous):

ok so it would be y=4x-7?

OpenStudy (anonymous):

why 4x?

OpenStudy (anonymous):

omg i mean positive 3x lol

OpenStudy (anonymous):

perfect :D so now you have the equation y=3x-7 and you know that the form is y=mx+b where m = slope and b= y intercept.. so.. whats the slope, and y intercept?

OpenStudy (anonymous):

the slope would be 3x and -7 would be the y intercept right?

OpenStudy (anonymous):

almost! since its mx+b, the slope is JUST the number before the x, which is 3 but you got the y intercept correct ^.^

OpenStudy (anonymous):

so now that you know the slope, you want ot take your point, (4,3) and plug it into the slope-intercept form which s Y-y1=m(x-x1) where your point is (x1,y1) and m = your slope

OpenStudy (anonymous):

go ahead and try it, and see what you come up with

OpenStudy (anonymous):

ok wait the m=3x?

OpenStudy (anonymous):

m is just 3

OpenStudy (anonymous):

x is just x

OpenStudy (anonymous):

oo ok wait let me c wat i get

OpenStudy (anonymous):

ok so i got 3x+3y=-3 am i right?

OpenStudy (anonymous):

not quite, remember you're only plugging your 4 and 3 where it says x1 and y1

OpenStudy (anonymous):

so in this case, x1=4, and y1=3 y-y1=m(x-x1)

OpenStudy (anonymous):

wana try one more time?

OpenStudy (anonymous):

yea sure thnx hold on

OpenStudy (anonymous):

so would it be x+3y=14?

OpenStudy (anonymous):

you want to get the y alone, here i'll show you y-y1=m(x-x1) y-3=3(x-4), i distributed the 3 y-3 = 3x -12, and here, you want to get the y alone so you move the 3 over by adding 3 to both sides y=3x-9

OpenStudy (anonymous):

that is the equation of the line that has a slope that is parallel to the first equation, and that passes through the points (4,3) ^.^ does that process make sense? do you want to recap?

OpenStudy (anonymous):

ooo ok it makes sense now thank you sooooo much!!!!

OpenStudy (anonymous):

but i have a question

OpenStudy (anonymous):

so first, you want to find the slope, and since that first line is parallel to your desired line, you know that the slope will be equivalent... you want to get it into the y=mx+b form though to know what m is then, using point slope form, the two points you were given, and the slope you found, you find the new equation of the line y-y1=m(x-x1)

OpenStudy (anonymous):

how dpo u find the slope cuz i look at the textbook and it wont give me a example can u plz expalin? so i could lean how to?

OpenStudy (anonymous):

sure, there are a few ways to do it.. so if you know that y=mx+b, and m = slope any time you see an equation of that form, you'll know what the slope is.. so like y=5x+7, slope = 5 y=2x+3, slope = 2

OpenStudy (anonymous):

two parallel lines will have the same slope, so if you know the slope of one, you'll know the slope of the other, and you can just identify it, again, by what number is before the x

OpenStudy (anonymous):

however, if you're not given an equation, but instead are given two points, there's a formula to find the slope... \[(y2-y1)/(x2-x1)\] where your two points are: (x1,y1) (x2, y2)

OpenStudy (anonymous):

ok cuz in my textbook there an example x-3y=6 and (-1,3) and the slope is 1/3 how is that possible?

OpenStudy (anonymous):

so for example, you're given the two points (1, 2) and (3,4) (x1, y1) (x1, y2) and you just plug them into the equation 4-2 / 3 -1 to get 2/2 or = 1

OpenStudy (anonymous):

so in this case, you want to again get the y by itself.. do you remember how to do that?

OpenStudy (anonymous):

u move everything to the other side?

OpenStudy (anonymous):

yup :o so walk me through the steps, what would i do first?

OpenStudy (anonymous):

this one is more tricky but i guess u would y alone and put on the other side 3?

OpenStudy (anonymous):

wait, run that by me one more time lol

OpenStudy (anonymous):

huh? lol

OpenStudy (anonymous):

i'm not sure what you just said haha, what would be my first step, if i was trying to solve this?

OpenStudy (anonymous):

u would move 3 on the other side to get y alone?

OpenStudy (anonymous):

yeah, but how would you move it?

OpenStudy (anonymous):

and that's actually not the first thing you want to do - i'll explain why in a second

OpenStudy (anonymous):

yea i have no clue cuz it doesnt make sence how the slope would be 1/3 and ok lol

OpenStudy (anonymous):

it actually is! i'll explain so you have x-3y=6, you said that you want to move the 3 over, the only way to move the 3 over would be to divide by 3, right? because to move things over, you want to do the opposite to what's being done to them, and in this case 3 is being multiplied by y, however, you can't just divide the -3y by 3, because the x is there too, you have to divide that by 3 as well.. so my suggestion would be to move the x over first.. how would you do that?

OpenStudy (anonymous):

ok so if i divide -3 with 6 it would be -2 right?

OpenStudy (anonymous):

yeah, it will, but you still have that x on the left side, and if you divide by -3, you have to divide that entire side by -3, not just the y.. so you can do it that way if you want, this is what you'll get x/-3 + y = -2

OpenStudy (anonymous):

so now what do you do to get it into y = mx+b?

OpenStudy (anonymous):

so it would be y=1/3x -2?

OpenStudy (anonymous):

yup!!! ^.^

OpenStudy (anonymous):

you understand why though right, you're not just looking at the book? :(

OpenStudy (anonymous):

well how do u get the 1 over the 3 wouldnt it be just 3?

OpenStudy (anonymous):

thats why i said it'd be easier if you moved the x over first x-3y=6, subtract the x from both sides you get -3y = -x +6 then you divide by -3 to get the y alone, and whatever you do to one side you have to do to the other y = (-x + 6)/-3 or y = -x/-3 + 6/-3

OpenStudy (anonymous):

y= x/3 -2 (1/3)(x) is equivalent to x/3 because you're just multiplying the x in and out of that 1

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

omg yes it does lol thank you but can u help me with my other problem?

OpenStudy (anonymous):

go ahead and close this one and ask again cuz this is starting to get super laggy

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

for sure ^.^

OpenStudy (anonymous):

wait let me just copy everything plz

OpenStudy (anonymous):

just the first 1 problem

OpenStudy (anonymous):

take your time haha no one's answering my questions so i'll be here for a while

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

you can actually close this one, and start a new one, and still look at this one while i answer the new one haha

OpenStudy (anonymous):

ok i have a question about the first problem

OpenStudy (anonymous):

o wait nvm got it lol

OpenStudy (anonymous):

ok im done ima close this 1 and start a new one